A 1000 kg weather rocket is launched straight up.The rocket motor provides a constant acceleration for 16 s, then the motor stops. The rocket altitude 20 s after launch is 6200 m. You can ignore any effects of air resistance. What was the rocket's acceleration during the first 16s? What is the rocket's speed as it passes through acloud 6200 m above the ground?
Answer #1


The rocket's altitude and velocity 16 s after the launch:
h1 = (1/2)a t
Answer #2

a).at t=0 rocket starts from rest
v=16a
distance s=0.5a(16)^2=128a
and given is
6200-128a=16a*(20-16)
6200=128a+64a
a=32.292 m/s^2

B).rocket speed is 16a at 6200 m
v=16*32.292=516.67 m/s